### v1

# TS-Haystack: A Multi-Scale Retrieval Benchmark for Time Series Language Models

Abstract
[…]
1 Introduction
Edge devices generate vast continuous streams of time series data: a single 100Hz accelerometer produces over 8 million datapoints daily. However, extracting insights typically relies on statistical pipelines or specialized models (Bulling et al., 2014). Time Series Language Models (TSLMs) provide a path toward reasoning over sensor data by bridging the gap between high-dimensional continuous signals and the embedding space of LLMs (Langer and others, 2025; Jin et al., 2024; Wang et al., 2025) using specialized time series encoders to bridge the dimensionality gap. However, these models are rarely trained on sample windows exceeding 1,000 datapoints, orders of magnitude below real-world requirements. We hypothesize that existing TSLM encoders do not retain sufficient temporal granularity at longer contexts lengths, and their compression mechanisms discard information needed for retrieval tasks.
Existing TSLM architectures face a fundamental trade-off between precise temporal representations and scalability. While full self-attention encoders attend to the whole sequence’s information, it incurs in a quadratic computational cost (Nie et al., 2023; Vaswani et al., 2017). In contrast, fixed latent-space architectures improve efficiency through compression, but at the expense of temporal resolution. We also hypothesize a task-dependent effect. In classification, latent resampling may benefit by filtering noise; for retrieval tasks requiring localization of specific events, the same compression may marginalize away distinguishing patterns. Although benchmarks probing long-context retrieval exist for LLMs (Kuratov et al., 2024; Kamradt, 2023; Wang et al., 2024) and recent efforts have enabled evaluating time series foundation models in classification and forecasting tasks (Goswami et al., 2024) and statistical understanding (Cai et al., 2024), to our best knowledge, no benchmark currently enables evaluation of retrieval performance over long-contexts for TSLMs.
To fill this gap, we introduce TS-Haystack, a retrieval benchmark adapting the needle-in-a-haystack paradigm to time series using real activity samples: short activity bouts (needles) are inserted into real accelerometer recordings spanning up to 2 hours (haystacks), where models must retrieve and reason about them. The benchmark comprises ten tasks with distinct requirements, enabling fine-grained diagnosis of where and why temporal understanding breaks down as context grows. We also evaluate different TSLM architectures on the Capture24 (Chan et al., 2024) classification dataset at context lengths ranging 2.56 seconds to 15 minutes. Further ablation introducing an oracle model variant allows us to investigate the origin of our long-context retrieval benchmark performance.
Figure 1: TS-Haystack construction pipeline. Left: A 1 hour window is sampled from a participant, activity needles ranging 1s – 6s are sampled at random. Right: Question-Answer pairs and chain-of-thought rationales are generated from plots and metadata-enriched prompt templates.
Contributions.
[…]
2 Background
We categorize TSLM architectures based on how they encode time series into the language model’s input space: full-resolution adapters that apply self-attention over all whole sequence length, and latent resampling methods, which compress them into a fixed number of tokens.
Full-Resolution Adapters.
Approaches such as Softprompting or Time-LLM (Jin et al., 2024), map the sequence of $N$ patch embeddings to the language space. However, this fidelity comes at a prohibitive computational cost ($\mathcal{O}(N^{2})$) either through the LLM’s extended context window or within the time series encoder, which must perform self-attention over the full patch sequence before projection (Wang et al., 2025). Consequently, these methods struggle to scale to larger context lengths due to computational cost and large memory requirements.
Latent Resampling.
[…]
3 Methods
In this section, we describe how we created the classification and TS-Haystack benchmark from the Capture24 dataset, including task taxonomy, dataset construction pipeline, needle insertion protocol, and training setup for the two evaluated architectures.
Table 1: Overview of TS-Haystack tasks grouped by category, with increasing difficulty.
Task Description Task Requirement Answer
Direct Retrieval
Existence Detect presence of a specific activity Pattern recognition Boolean
Localization Temporally ground an activity Recognition, Localization Time range
Counting Enumerate activity occurrences Recognition, Enumeration Integer
Temporal Reasoning
Ordering Compare activity temporal order Localization, comparison Boolean
State Query Identify regime around local event Cross-scale integration Category
Antecedent Identify activity preceding a target Localization, Adjacency Category
Multi-Step Reasoning
Comparison Compare durations across gaps Localization, comparison Time range
Multi-Hop Locate $K$-th target relative to anchor Localization, counting Time range
Contextual Anomaly
Anomaly Det. Detect cross-regime activity Contextual reasoning Bool. + cat.
Anomaly Loc. Detect and localize anomaly Contextual, Localization Bool. + cat. + time
Note: Det. - Detection; Loc. - Localization
Classification on Capture24.
We first evaluate both architectures on activity classification using Capture24 (Chan et al., 2024), a dataset of 151 participants wearing triaxial wrist accelerometers at 100 Hz for approximately 24 hours. We use the Willetts et al. (2018) annotation scheme as the class mapping for the annotated Capture24 dataset, providing 10 activity classes. Full details on class distributions, sampling budgets, and training protocol are provided in Appendix A.
TS-Haystack Benchmark.
Following the needle-in-a-haystack paradigm (Kamradt, 2023), we insert short activity bouts (needles) into longer background recordings (haystacks) and task the model with retrieving and reasoning about them. The benchmark comprises ten tasks of increasing complexity (Table 1), generated at six context lengths. For each combination of context length and task type, we generate 1,000 training, 150 validation, and 150 test samples from 20 manually crafted question templates per task. Full task specifications and examples are provided in Appendices B.7 and B.
Needle Insertion.
Needles are randomly sampled and inserted using per-channel mean alignment to adjust for participant-specific sensor biases while preserving activity dynamics, with cosine blending at insertion boundaries for smooth transitions (Appendix B.2). We validate that this protocol produces statistically undetectable insertions via classifier-based testing (Appendix B.3).
Training Details.
We select recent open-source implementations of the two architectural categories: OpenTSLM-Flamingo (Langer and others, 2025) and ITFormer (Wang et al., 2025), using their default LLM backbones: Llama-3.2-1B (Grattafiori and others, 2024) and Qwen2.5-1.5B-Instruct (Qwen Team, 2025). Models are trained jointly on all tasks and context lengths, using chain-of-thought rationales generated with gpt-4.1-mini-2025-04-14 (OpenAI, 2023). ITFormer encountered out-of-memory errors at context lengths $\geq$40,000 datapoints; these were excluded from training and evaluation (Figure 2).
Oracle Variant.
Since both the time series encoder and the language model contribute to the final answer, poor retrieval performance could stem from either component. We train an Oracle Variant of OpenTSLM-Flamingo receiving ground-truth activity segmentation as text metadata alongside the time series input, bypassing the need for the encoder to provide grounded activity recognition in its output representations. If the oracle maintains high retrieval accuracy across context lengths, the bottleneck lies in the time series encoding rather than in reasoning difficulty (see Appendix B.6).
4 Results
In the following, we assess scaling behavior on classification and TS-Haystack retrieval tasks.
Scaling Behavior.
[…]
Oracle Upper Bound.
Our oracle variant achieves 86.8% overall accuracy and, critically, maintains stable performance across context lengths (82.2%–91.3%), indicating that the OpenTSLM-Flamingo LLM backbone possesses sufficient reasoning capacity.
(a) Classification (Capture24)
(b) Retrieval (TS-Haystack)
Figure 2: Scaling behaviors. (a) Capture24 classification: Macro-F1 and Balanced Accuracy increase with context length. (b) TS-Haystack: Retrieval accuracy degrades with context length.
Table 2: TS-Haystack accuracy by task and context length. ITFormer encounters OOM errors beyond 100s; long-context results are OpenTSLM-Flamingo only. Best results per time series length in bold.
2.56s 10s 100s 15min 1h 2h
Task Rand % Flam. IT Flam. IT Flam. IT Flam. Flam. Flam.
Direct Retrieval
Existence 50.0 57.3 40.0 47.3 48.0 54.0 42.7 53.3 51.3 50.0
Localization $\scriptstyle\sim$1.5 2.7 2.0 2.0 2.7 6.0 1.3 4.0 3.3 2.7
Counting 20.0 19.3 12.7 17.3 17.3 20.7 20.0 16.0 20.7 18.7
Temporal Reasoning
Ordering 50.0 46.7 30.7 52.7 35.3 38.7 30.0 41.3 54.0 50.7
State Query 10.0 65.3 4.0 68.0 3.3 24.7 4.7 22.0 31.3 34.7
Antecedent 10.0 7.3 7.3 13.3 6.0 11.3 5.3 8.7 9.3 9.3
Multi-Step Reasoning
Comparison $\scriptstyle\sim$1.5 8.7 2.0 9.3 0.7 6.7 1.3 6.0 8.0 3.3
Multi-Hop $\scriptstyle\sim$1.5 2.7 0.0 0.7 1.3 2.7 0.0 3.3 1.3 2.7
Contextual Anomaly
Anomaly Det. 50.0 52.0 52.0 46.7 52.7 45.3 48.7 50.0 48.0 54.7
Anomaly Loc. $\scriptstyle\sim$0.75 34.0 0.7 30.7 0.0 34.0 0.0 27.3 28.7 24.0
Average $\scriptstyle\sim$19.5 29.6 15.1 28.8 16.7 24.4 15.4 23.2 25.6 25.1
Discussion.
With our work, we seek to improve understanding of TSLM architectures with respect to task type and context length, expanding existing benchmarks from annotated short intervals (Goswami et al., 2024) and synthetic long contexts (Cai et al., 2024) to labeled long-context retrieval.
Our results confirm that TSLMs experience task-dependent performance: OpenTSLM-Flamingo improves in classification as context length grows, yet retrieval performance degrades over the same range. Contrary to our initial hypothesis, retrieval performance does not linearly degrade with context length; performance drops from 29.6% (2.56 s) to 23.2% (15 min) then plateaus near 25% through increasing context lengths up to 2 hours. This non-monotonic degradation suggests that the bottleneck is not purely one of information loss through compression. State Query performance drops from 65.3% at 2.56 s to 24.7% at 15 min, contributing to 78% of the OpenTSLM-Flamingo degradation over the same period. At short lengths, State Query resembles a classification problem: the background is dominated by one activity regime that frequently coincides with the needle’s context, making correct answers achievable through background classification. Other tasks requiring localization or contextualized short-interval identification remain difficult across all context lengths, suggesting a shared bottleneck even at short sequences. We hypothesize that training on all context lengths jointly may have induced the model to converge toward a context-invariant strategy, learning to retain global features rather than localized patterns irrespective of the true compression ratio, motivating future work on curriculum learning. Our oracle experiment suggests that this degradation originates within the time series encoding and projection layers: when ground-truth segmentation is provided as text, retrieval accuracy remains stable across all context lengths (82.2%–91.3%), indicating sufficient reasoning capacity in the LLM backbone. All findings point to a key design principle: temporal granularity must be decoupled from computational complexity throughout the architecture, not only at the adapter interface.
We note that ITFormer, despite using full-attention, underperforms even at short contexts, falling below random baseline in retrieval. An extended discussion is provided in Appendix B.5 and results should be interpreted with caution. However, the consistently flat accuracy suggests that preserving full-attention in the encoder is insufficient for retrieval.
We introduce TS-Haystack as a diagnostic benchmark for future research in long-context architectures (Gu and Dao, 2023; Behrouz et al., 2025), incorporating task and context-length adaptation.
Appendix A Classification Experiment Details
A.1 Experimental Setup
Model architecture.
We use the Flamingo variant of OpenTSLM (Langer and others, 2025) with Llama-3.2-1B as the backbone LLM. To accommodate variable context lengths beyond the original training configuration, we extend the positional encoding of the Perceiver Resampler via linear interpolation. Critically, these interpolated positional embeddings are unfrozen during fine-tuning, allowing the model to adapt to longer temporal contexts.
Dataset alignment.
We downsample Capture24 from 100Hz to 50Hz to match the sampling rate of the HAT accelerometer dataset used for OpenTSLM pretraining. The minimum context length of 2.56 seconds (128 timesteps) corresponds to the original OpenTSLM training window, providing a consistent baseline for measuring scaling behavior. This alignment ensures that performance differences across context lengths reflect genuine scaling effects rather than distribution shift from mismatched preprocessing.
Label assignment.
Since longer context windows may span multiple activities, we determine classification labels via majority voting over the ground-truth annotation timeline. We require at least 60% temporal coverage by a single activity class for a sample to receive that label; windows without a clear majority are excluded from evaluation.
Training protocol.
To ensure no data leakage, we first split participants (total of 151) into train (100), validation (25) and test (26) splits at random. We adopt a curriculum learning approach, training sequentially from shorter to longer contexts. For each context length, we use a maximum sample budget: 80,000 training, 15,000 validation, and 15,000 test samples (see Tables 3 and 4 for the relationship between available and sampled data). Models are trained for up to 50 epochs with early stopping based on validation loss, using a learning rate of $5\times 10^{-5}$ with AdamW optimizer.
Dataset size.
While longer context lengths up to 1 hour were initially included in training and yielded increasingly higher Macro-F1 scores, evaluation at these scales is limited by the scarcity of sufficiently long single-activity bouts in the source data; we therefore exclude these lengths from our main results to ensure statistically robust comparisons. At 5 and 15 minutes context length, the classification evaluation relies on smaller test sets (4,697 and 1,473 samples) due to the scarcity of contiguous single-activity segments. Performance at these context lengths should be interpreted with caution; yet the consistent upward trend suggests a robust positive effect.
A.2 Dataset Statistics
Table 3 reports the total number of qualifying activity bouts (i.e. windows meeting the 60% majority criterion) available in the Capture24 dataset for each context length and split. As context length increases, the number of non-overlapping windows that satisfy the majority-vote criterion drops sharply: the 2.56 s setting yields over 3.5 M bouts, whereas the 15 min setting provides fewer than 6,500.
Table 3: Total number of qualifying activity bouts in Capture24 per context length and data split, before any sampling.
Context Train Val Test Total
2.56 s 2,488,532 542,673 566,768 3,597,973
10 s 636,107 138,644 144,841 919,592
60 s 106,515 23,292 24,295 154,102
5 min 20,724 4,502 4,697 29,923
15 min 6,489 1,406 1,473 9,368
Avg. split 69.2% 15.1% 15.7% 100%
To maintain a consistent training signal across context lengths despite the large variation in available data, we cap sampling at 80,000 training, 15,000 validation, and 15,000 test samples per context length (Table 4). When the pool exceeds the cap, samples are drawn uniformly at random (seed 42); otherwise the full set is used. This strategy ensures that shorter-context models are not disproportionately advantaged by larger datasets, while longer-context settings retain every available sample.
Table 4: Actually sampled dataset sizes per context length and split. “Random” indicates uniform subsampling to the budget cap; “full” indicates all available bouts were used.
Train Val Test
Context NN Method NN Method NN Method
2.56 s 80,000 random 15,000 random 15,000 random
10 s 80,000 random 15,000 random 15,000 random
60 s 80,000 random 15,000 random 15,000 random
5 min 20,724 full 4,502 full 4,697 full
15 min 6,489 full 1,406 full 1,473 full
Tables 5– 7 report the per-class label distributions for the sampled training, validation, and test sets, respectively. Across all splits and context lengths, sleep and sitting together account for roughly 70% of samples, whereas minority classes such as bicycling, sports, and manual-work each contribute fewer than 2%. This imbalance is further exacerbated at longer contexts, where rare activities are increasingly unlikely to sustain the 60% majority threshold over extended windows.
Table 5: Training set class distribution (number of samples) across context lengths on Capture24. Percentages are computed within each context length.
Activity 2.56 s 10 s 60 s 5 min 15 min
sleep 29,538 (36.9%) 30,048 (37.6%) 29,481 (36.9%) 7,851 (37.9%) 2,601 (40.1%)
sitting 28,712 (35.9%) 28,450 (35.6%) 28,661 (35.8%) 7,637 (36.9%) 2,515 (38.8%)
household-chores 5,616 (7.0%) 5,460 (6.8%) 5,540 (6.9%) 1,383 (6.7%) 381 (5.9%)
walking 5,107 (6.4%) 4,944 (6.2%) 5,114 (6.4%) 1,080 (5.2%) 220 (3.4%)
mixed-activity 3,324 (4.2%) 3,422 (4.3%) 3,440 (4.3%) 869 (4.2%) 263 (4.1%)
vehicle 2,849 (3.6%) 2,883 (3.6%) 2,882 (3.6%) 747 (3.6%) 231 (3.6%)
standing 2,515 (3.1%) 2,485 (3.1%) 2,531 (3.2%) 559 (2.7%) 94 (1.4%)
manual-work 1,074 (1.3%) 1,063 (1.3%) 1,091 (1.4%) 281 (1.4%) 89 (1.4%)
bicycling 848 (1.1%) 839 (1.0%) 848 (1.1%) 209 (1.0%) 59 (0.9%)
sports 417 (0.5%) 406 (0.5%) 412 (0.5%) 108 (0.5%) 36 (0.6%)
Total 80,000 80,000 80,000 20,724 6,489
Table 6: Validation set class distribution (number of samples) across context lengths on Capture24. Percentages are computed within each context length.
Activity 2.56 s 10 s 60 s 5 min 15 min
sleep 5,545 (37.0%) 5,553 (37.0%) 5,502 (36.7%) 1,715 (38.1%) 569 (40.5%)
sitting 5,521 (36.8%) 5,462 (36.4%) 5,513 (36.8%) 1,695 (37.6%) 554 (39.4%)
walking 1,032 (6.9%) 1,047 (7.0%) 1,040 (6.9%) 268 (6.0%) 52 (3.7%)
household-chores 794 (5.3%) 761 (5.1%) 816 (5.4%) 222 (4.9%) 56 (4.0%)
mixed-activity 634 (4.2%) 634 (4.2%) 647 (4.3%) 185 (4.1%) 59 (4.2%)
standing 615 (4.1%) 632 (4.2%) 613 (4.1%) 153 (3.4%) 31 (2.2%)
vehicle 562 (3.7%) 633 (4.2%) 578 (3.9%) 176 (3.9%) 57 (4.1%)
bicycling 216 (1.4%) 197 (1.3%) 217 (1.4%) 67 (1.5%) 21 (1.5%)
sports 50 (0.3%) 49 (0.3%) 37 (0.2%) 13 (0.3%) 4 (0.3%)
manual-work 31 (0.2%) 32 (0.2%) 37 (0.2%) 8 (0.2%) 3 (0.2%)
Total 15,000 15,000 15,000 4,502 1,406
Table 7: Test set class distribution (number of samples) across context lengths on Capture24. Percentages are computed within each context length. Longer contexts yield fewer non-overlapping windows from the fixed-length recordings, reducing total test samples at 5 min and 15 min.
Activity 2.56 s 10 s 60 s 5 min 15 min
sleep 5,424 (36.2%) 5,522 (36.8%) 5,423 (36.2%) 1,758 (37.4%) 584 (39.6%)
sitting 5,183 (34.6%) 5,087 (33.9%) 5,186 (34.6%) 1,659 (35.3%) 555 (37.7%)
household-chores 1,379 (9.2%) 1,377 (9.2%) 1,404 (9.4%) 418 (8.9%) 117 (7.9%)
walking 1,041 (6.9%) 1,065 (7.1%) 1,061 (7.1%) 272 (5.8%) 48 (3.3%)
vehicle 659 (4.4%) 669 (4.5%) 621 (4.1%) 212 (4.5%) 64 (4.3%)
mixed-activity 566 (3.8%) 559 (3.7%) 580 (3.9%) 174 (3.7%) 55 (3.7%)
standing 407 (2.7%) 397 (2.6%) 419 (2.8%) 101 (2.2%) 22 (1.5%)
sports 171 (1.1%) 169 (1.1%) 164 (1.1%) 54 (1.1%) 16 (1.1%)
manual-work 139 (0.9%) 117 (0.8%) 116 (0.8%) 40 (0.9%) 11 (0.7%)
bicycling 31 (0.2%) 38 (0.3%) 26 (0.2%) 9 (0.2%) 1 (0.1%)
Total 15,000 15,000 15,000 4,697 1,473
A.3 Per-Class Performance
Class imbalance.
The Capture24 test set exhibits severe class imbalance: sleep and sitting together account for over 70% of all windows at every context length, while minority classes such as bicycling, manual-work, and standing each represent fewer than 3% of samples (Table 5). To mitigate that performance on rare activities is not masked by majority-class accuracy while preserving a significant number of samples for training and evaluation at longer context lengths, we report Macro-F1 and Balanced Accuracy as our primary evaluation metrics Bulling et al. (2014). Macro-F1 computes the unweighted mean of per-class F1 scores, giving equal importance to every activity regardless of prevalence. Balanced Accuracy averages per-class recall, providing a complementary view that is similarly insensitive to class frequency. These metrics penalize models that achieve high overall accuracy by defaulting to majority-class predictions, a pattern evident in the prediction distributions of Table 7, where the model over-predicts sleep and sitting across all context lengths.
Table 8 reports per-class F1 scores alongside Macro-F1 and Balanced Accuracy across context lengths.
Table 8: OpenTSLM-Flamingo per-class F1 scores (%) across context lengths on Capture24.
Activity 2.56 s 10 s 60 s 5 min 15 min
sleep 83.2 86.2 85.1 86.8 88.0
sitting 60.4 67.0 68.3 73.3 74.8
household-chores 26.2 38.7 42.1 45.7 57.7
vehicle 55.7 50.1 54.7 55.7 47.5
walking 32.0 37.6 36.4 41.3 34.7
bicycling 40.0 55.2 64.0 70.0 66.7
sports 24.9 27.4 26.6 19.4 31.6
mixed-activity 4.8 5.4 9.7 6.6 13.3
standing 0.0 0.0 0.0 0.0 0.0
manual-work 1.3 0.0 2.8 2.9 0.0
Macro-F1 32.9 36.8 39.0 40.2 41.4
Balanced Accuracy 33.1 35.7 39.8 39.4 42.4
A.4 Training Statistics
Table 9 compares class-imbalance-aware metrics across context lengths for OpenTSLM-Flamingo and ITFormer.
Table 9: Evaluation metrics across context lengths. At 15 min, ITFormer encountered out-of-memory errors (†).
OpenTSLM-Flamingo ITFormer
Duration Timesteps Macro-F1 Bal. Acc. Macro-F1 Bal. Acc.
2.56 s 128 32.9% 33.1% 13.9% 17.8%
10 s 500 36.8% 35.7% 21.8% 25.4%
60 s 3,000 39.0% 39.8% 21.6% 25.0%
5 min 15,000 40.2% 39.4% 15.8% 17.3%
15 min 45,000 41.4% 42.4% OOM
OpenTSLM-Flamingo compression ratio.
[…]
ITFormer baseline.
We additionally evaluate ITFormer (Wang et al., 2025) as a full-resolution baseline. We adopt the original architecture with a 4-layer time series encoder ($d_{\text{model}}{=}512$, 8 heads) and a 2-layer cross-modal prefix module ($d_{\text{model}}{=}896$, 16 heads, 25 prefix tokens) projecting into a Qwen2.5-1.5B-Instruct backbone. Patches use a non-overlapping window of 60 samples at 50 Hz (1.2 s per patch). We train with a learning rate of $1\times 10^{-5}$, batch size 8, weight decay $1\times 10^{-6}$, and FP16 mixed precision for 3 epochs. The LLM backbone is frozen and only the time series encoder and prefix module are trained. At 15 minutes (900 s context), training encountered out-of-memory errors on an NVIDIA RTX PRO 6000 (96 GB).
Appendix B TS-Haystack details
B.1 Related Benchmark Literature
TS-Haystack builds upon and distinguishes itself from prior work in time series question answering and long-context evaluation. We position our benchmark relative to existing approaches across three dimensions: task design philosophy, data construction methodology, and evaluation scope.
Time Series Foundation Models.
MOMENT (Goswami et al., 2024) releases the first open-source foundation models for general-purpose time series analysis, pre-trained on the Time Series Pile, a large multi-domain collection of public time series. Their benchmark evaluates five tasks: forecasting, classification, anomaly detection, and imputation, with emphasis on limited supervision settings. However, as noted by the authors, most datasets found operate on a fixed-length windows of T=512T{=}512 timesteps and focus on holistic tasks that do not require localizing or reasoning about specific events within extended temporal contexts. TS-Haystack complements this line of work, testing whether models preserve fine-grained temporal information across contexts spanning orders of magnitude beyond typical training windows.
Time Series Understanding Benchmarks.
TimeSeriesExam (Cai et al., 2024) evaluates LLMs across five core time series understanding categories: pattern recognition, noise understanding, similarity analysis, anomaly detection, and causality analysis. While their procedural generation approach enables scalability, the benchmark relies on synthetic time series and focuses primarily on statistical properties (e.g., “Do the two time series have similar variance?”, “What is the trend?”). In contrast, TS-Haystack emphasizes semantic interpretation using real sensor data with activity labels, testing whether models can derive real-world meaning from temporal patterns rather than merely describing statistical features.
Time-MQA (Kong et al., 2025) is more closely aligned with our objective, unifying forecasting, imputation, anomaly detection, classification, and open-ended reasoning under a question-answering framework. Their TSQA dataset comprises approximately 200k question-answer pairs across diverse domains. However, Time-MQA does not specifically probe long-context retrieval, the ability to locate and reason about specific events within extended temporal contexts.
Long-Context Evaluation in Other Modalities.
The needle-in-a-haystack paradigm originated in LLM evaluation (Kamradt, 2023), where a specific fact is embedded within long text documents to test retrieval accuracy. BABILong (Kuratov et al., 2024) extends this to reasoning tasks, demonstrating that popular LLMs effectively utilize only 10–20% of their claimed context length. For vision-language models, MMNeedle (Wang et al., 2024) adapts the paradigm to multimodal retrieval, stress-testing models’ ability to locate target sub-images within large visual contexts. To our knowledge, TS-Haystack is the first benchmark to apply the needle-in-a-haystack methodology to continuous time series data, where the challenge is compounded by the absence of discrete token boundaries and the need to preserve temporal precision under aggressive compression.
Distinguishing Contributions.
Unlike synthetic benchmarks (Cai et al., 2024), TS-Haystack uses real accelerometer recordings with ecologically valid activity patterns. Our statistically validated needle insertion protocol (Appendix B.3) ensures that inserted activities are indistinguishable from natural occurrences, enabling scalable semi-synthetic dataset generation without introducing detectable artifacts. This methodology may prove valuable beyond benchmarking, for instance, in augmenting training data for domains where anomalies are scarce, such as rare event detection in healthcare or industrial monitoring.
B.2 Dataset Construction
Pre-processing.
For each participant in Capture24, we construct an activity timeline by merging consecutive annotations of the same activity into bouts, recording start time, end time, and duration. We aggregate all bouts across participants into a bout index organized by activity class, enabling efficient sampling of needle candidates by activity type and duration.
Needle selection.
We deliberately avoid sampling according to empirical transition probabilities: real activity sequences exhibit strong temporal dependencies that standard HAR pipelines exploit via HMM smoothing (Chan et al., 2024). Uniform random needle-background pairings ensure models cannot game the benchmark through learned Markovian priors, requiring genuine temporal retrieval for correct answers.
Needle insertion.
To ensure inserted needles are statistically consistent with the surrounding context, we apply per-channel mean alignment:
$\mathbf{x}^{\prime}=\mathbf{x}_{\text{needle}}-\bm{\mu}_{\text{needle}}+\bm{\mu}_{\text{target}}$ (1)
where $\bm{\mu}$ denotes the per-channel mean. This adjusts for participant-specific sensor biases (orientation, placement) while preserving the activity’s characteristic amplitude and temporal dynamics. To ensure smooth transitions at insertion boundaries, we apply cosine blending over a variable window of $w$ samples, weighting the needle contribution as $\alpha(t)=\tfrac{1}{2}(1-\cos(\pi t/w))$ to gradually interpolate between background and needle signals.
Question generation.
Questions are generated from 20 manually-crafted templates per task, with activity names and timestamps filled programmatically per sample. Contextual timestamps are included in the input prompts and grounded in real clock time (e.g., “02:34:56:789 AM”) derived from the original Capture24 recording metadata, enabling naturalistic queries aligned with real-world applications. All experiments use a fixed master seed with deterministic sample generation for reproducibility.
Configurable difficulty.
TS-Haystack supports configurable difficulty along multiple axes: needle position (beginning/middle/end/random), needle duration, minimum gap between bouts, distractor density, and other task-specific parameters. For our main dataset generation, we use random needle positions with the task-specific settings shown in Table 10.
Table 10: Task configuration parameters for TS-Haystack generation. Needle length is specified as a percentage of context length. All tasks use 2% margin and minimum gap ratios (capped at 100 samples for longer contexts).
Task Needle Length Background Task-Specific
Existence 2–10% any 1 – 3 distractors
Localization 2–10% any 1 – 3 distractors
Counting 2–8% any 1 – 5 bouts
Ordering 2–10% any
State Query 1–5% mixed 2 – 5 states, min 20% duration
Antecedent 2–8% any 10-sample adjacency gap
Comparison 2–8% any 2 – 4 bouts, min 2% duration diff
Multi-Hop 2–6% any $K$ position index[$\sim[0.4,0.4,0.2]$
Anomaly Detection 3–15% pure 1 – 3 distractors
Anomaly Localization 3–15% pure 1 – 3 distractors
Chain-of-thought annotations.
For each sample, we generate chain-of-thought (CoT) rationales using ChatGPT 4.1 mini (OpenAI, 2023), conditioned on image plots of the accelerometer data along with rich metadata including activity timelines, bout boundaries, and signal statistics. These annotations support training and evaluation of reasoning capabilities beyond direct answer extraction.
B.3 Insertion Quality Validation
To verify that our style transfer and cosine blending protocol produces clean insertions without detectable artifacts, we design a classifier-based validation test. The key insight is that if a machine learning model cannot distinguish samples with insertions from samples without, the insertion process introduces no systematic artifacts.
Methodology.
We construct a binary classification task:
  • Class 0 (negative): Pure background windows with no insertion.
  • Class 1 (positive): Pure background windows with a same-activity needle inserted from a different participant.
Using same-activity needles isolates insertion artifacts from activity-class differences. If the classifier succeeds, it must be detecting blending imperfections rather than activity signatures.
We generate 5,000 training samples and 500 test samples at 2 and 3 seconds context length (200 and 300 samples at 100Hz), with needle lengths ranging from 2–8% of context aligned with the true dataset configurations and balanced class distributions. An XGBoost classifier (Chen and Guestrin, 2016) is trained on flattened triaxial accelerometer features (600 and 900 dimensions) using 100 estimators with max depth 6.
Results.
The classifier achieves an AUC of 0.499 and 0.490 respectively for different context lengths, close to random chance (0.50). This confirms that our mean-shift normalization combined with cosine blending produces insertions that are statistically undetectable, even to a gradient-boosted ensemble operating on raw signal features. The protocol successfully preserves activity dynamics while eliminating participant-specific biases and boundary discontinuities.
B.4 ITFormer Extended Discussion
Several limitations of our ITFormer evaluation warrant discussion. First, we did not perform hyperparameter tuning specific to the Capture24 task or TS-Haystack tasks; the learning rate, batch size, and training duration were adopted from the original work and may be suboptimal for accelerometer-based activity recognition. Second, the LLM backbone was kept frozen: finetuning or using LoRA adapters could improve the model’s ability to interpret time series prefix embeddings. Finally, OOM errors limited the model from learning representations from longer context lenghts, effectively training on a smaller dataset than OpenTSLM-Flamingo. Further ablation experiments on patching strategies could alleviate memory constraints or preserve timeseries features of the specific dataset. Future work could investigate these directions to determine whether the performance gap with latent resampling methods narrows under a more optimized full-resolution setup.
B.5 ITFormer Extended Discussion
Our ITFormer evaluation adopted hyperparameters directly from the original work without task-specific tuning, and kept the LLM backbone frozen, while ablating on the time series encoder unfreezing, which resulted in marginal improvements used in the final results section. Both choices may disadvantage ITFormer relative to OpenTSLM-Flamingo, where the backbone is finetuned. This suggests that the retrieval bottleneck is not solely a matter of compression ratio: architectural choices such as encoder depth, patch granularity, frozen versus tuned LLM backbones, and the choice of backbone itself (Qwen2.5-1.5B-Instruct vs. Llama-3.2-1B) mediate whether temporal detail is preserved in a form the language model can exploit. Additionally, alternative patching strategies could alleviate memory constraints at longer contexts. A controlled comparison isolating these factors remains an important direction for future work.
B.6 Oracle Model Experiment
To further isolate time series representation as the bottleneck for TSLM performance on our benchmark, we introduce an Oracle-based upper bound evaluation. The oracle model receives the ground-truth activity timeline as text prepended to the standard prompt as shown in Table 11, bypassing the requirement for the time series encoder to provide the activity recognition and temporal grounding information in its representations for successful answering.
Table 11: Oracle prompt for an example multi-hop task. Italic text indicates the oracle-only addition; regular text is shared across configurations.
Prompt (Pre-prompt Oracle + Regular)
Activity Timeline (Ground Truth):
Recording: 10:53:45 PM – 10:53:55 PM
10:53:45 – 10:53:50: sleep
10:53:50 – 10:53:50: sitting [inserted]
10:53:50 – 10:53:51: sleep
10:53:51 – 10:53:51: sitting [inserted]
10:53:51 – 10:53:51: sleep
10:53:51 – 10:53:52: mixed-act. [inserted]
10:53:52 – 10:53:55: sleep
You are given accelerometer data in all three dimensions from a wrist-worn sensor. The recording spans from 10:53:45 PM to 10:53:55 PM.
Question: When was the second sitting activity before mixed-activity recorded?
[accelerometer data x/y/z axes]
Instructions: Analyze the accelerometer data carefully. Think step-by-step about what the signal patterns indicate. Write your reasoning as a natural paragraph. End your response with “Answer: <your answer>”
B.7 Task Details
In the following, we provide detailed descriptions with visual examples sampled from our sample generation pipeline for each task. Task variations such as non Existence, non Anomaly Detection or all 8 Comparison variants are excluded for brevity.
Task 1: Existence.
“Is there {activity} in this recording?”
This task tests whether the model can detect the presence of a specific activity pattern within a longer recording. To prevent variance-based detection shortcuts (where models detect any insertion rather than the specific activity), we insert multiple same-regime distractors. The model must learn activity-specific patterns rather than relying on statistical anomalies.
Construction: (1) Sample background window; (2) Select activity regime (sedentary/active); (3) Insert $N$ needles from regime activities not in background; (4) For positive samples, ask about an inserted activity; for negative, ask about a non-inserted activity from the same regime.
Example: Q: “Is there manual-work in this recording?” A: “No.”
Figure 3: Task 1: Existence. The model must detect whether a specific activity is present among same-regime distractors.
Task 2: Localization.
“When did the {activity} bout occur?”
Extends existence by requiring temporal grounding. The model must not only detect the activity but report its precise time range. Distractors with similar statistical properties force the model to identify activity-specific temporal signatures.
Construction: Same as Existence, but answer requires the timestamp range of the target needle.
Example: Q: “When did the sports bout occur?” A: “From 02:09:52:635 to 02:10:08:967.”
Figure 4: Task 2 (Localization). The model must identify and temporally ground a specific activity
Task 3: Counting.
“How many {activity} bouts occurred?”
Tests enumeration capabilities: the model must detect all instances of an activity and count them correctly. Minimum gaps between bouts ensure distinguishability while making the task non-trivial.
Construction: (1) Sample background; (2) Sample $N\in[1,5]$ bouts of target activity; (3) Insert at non-overlapping positions with minimum gap constraint.
Example: Q: “How many standing bouts occurred?” A: “3.”
Figure 5: Task 3 (Counting). The model must enumerate all occurrences of a specific activity
Task 4: Ordering.
“Did {activity_a} occur before {activity_b}?”
Tests temporal comparison between two distinct activities. We use needle insertion rather than natural bout selection to prevent models from exploiting learned activity transition statistics. Using random needle sampling pairings instead of existing activity transitions in the data ensure models cannot game the benchmark through learned Markovian priors, requiring genuine temporal retrieval for correct answers (Chan et al., 2024).
Construction: (1) Sample two distinct activities $A$ and $B$; (2) Sample background excluding both; (3) Randomly assign temporal order (50/50 balance); (4) Insert both needles sequentially.
Example: Q: “Did sports occur before sleep?” A: “Yes.”
Figure 6: Task 4 (Ordering). The model must determine the temporal relationship between two activities
Task 5: State Query.
“What was the activity level when {event} occurred?”
Tests cross-scale integration: the model must detect a local event (needle, seconds-scale) and identify the surrounding global activity regime (minutes-scale). This requires understanding hierarchical temporal context.
Construction: (1) Sample mixed background with 2–5 activity states; (2) Select a global state region; (3) Insert needle within that region; (4) Ask about the global state, not the needle activity.
Example: Q: “What was the overall activity when the vehicle activity occurred?” A: “Standing.”
Figure 7: Task 5 (State Query). The model must identify the global activity regime surrounding a local event
Task 6: Antecedent.
“What activity occurred immediately before {target}?”
Tests sequential reasoning: the model must locate the target activity and identify what directly preceded it. Two-needle insertion gives control over the antecedent-target pairing, preventing reliance on natural transition statistics.
Construction: (1) Sample background (prefer low-activity); (2) Sample antecedent $A$ and target $T$; (3) Insert $A$ followed immediately by $T$ with small adjacency gap.
Example: Q: “Looking at the sequence, what came before bicycling?” A: “Sports.”
Figure 8: Task 6 (Antecedent). The model must identify the activity immediately preceding a target
Task 7: Comparison.
“What was the {longest/shortest} period {with/without} {activity}?”
Tests comparison reasoning with four question variants: longest/shortest $\times$ with/without. The “without” polarity requires understanding negation and computing gaps between activity bouts. All inserted bouts have distinct durations to ensure unambiguous answers.
Construction: (1) Sample question type (extremum $\times$ polarity); (2) Insert $N\geq 2$ bouts with distinct durations; (3) For “with”: answer is extremum bout; for “without”: answer is extremum gap.
Example: Q: “What was the longest period without household-chores?” A: “From 8:32:52:619 to 08:35:30:682.”
Figure 9: Task 7 (Comparison). The model must compare durations across multiple activity bouts or gaps
Task 8: Multi-Hop.
“When did the K-th {target} occur {before/after} {anchor}?”
Tests multi-step reasoning: (1) locate the anchor activity, (2) identify target bouts relative to anchor, (3) count to the $K$-th occurrence in the specified direction. Optional distractors on the opposite side test directional understanding.
Construction: (1) Sample anchor $A$ and target $T$; (2) Sample $K\in\{1,2,3\}$ and direction; (3) Position anchor with room for $K$ targets; (4) Insert $K$ target needles in order.
Example: Q: “When did the third sports bout occur after mixed-activity?” A: “From 09:38:32:874 to 09:38:40:924.”
Figure 10: Task 8 (Multi-Hop). The model must locate an anchor, then count to the $K$-th target in a specified direction
Task 9: Anomaly Detection.
“Is there an anomaly in this recording?”
Tests contextual reasoning: the model must detect activity that is anomalous relative to the background regime. Unlike Existence (where the target is given), the model must identify what constitutes an anomaly. We define anomalies as cross-regime insertions (e.g., vigorous activity in a sedentary background). We partition activities into two regimes: sedentary (sleep, sitting, standing, vehicle) and active (walking, mixed-activity, bicycling, manual-work, sports, household-chores). Mandatory same-regime distractors prevent “insertion detected” shortcuts: both positive and negative samples contain insertions, so the model must identify regime mismatch.
Construction: (1) Sample pure background from one regime; (2) Positive: insert 1 cross-regime needle (anomaly) + $N$ same-regime distractors; Negative: insert only same-regime distractors.
Example: Q: “Is there an anomaly in this recording?” A: “Yes, there is anomalous sitting activity in the active background.”
Figure 11: Task 9 (Anomaly Detection). The model must identify cross-regime activity without being told what to look for
Task 10: Anomaly Localization.
“Is there an anomaly, and if so, when did it occur?”
Extends anomaly detection by requiring temporal localization. The model must detect the cross-regime violation and specify its time range. This combines contextual reasoning with precise temporal grounding.
Construction: Same as Anomaly Detection, but positive answers must include the time range of the anomalous activity.
Example: Q: “Is there an anomaly, and if so, when?” A: “Yes, there is anomalous sitting activity from 04:21:20:792 to 04:21:34:333.”
Figure 12: Task 10 (Anomaly Localization). The model must detect and temporally ground cross-regime anomalies